OpenStudy (cydney_morgan):

Elements in the same group on the periodic table have the same number of electrons. number of protons. <--- My choice number of neutrons. number of valence electrons.

OpenStudy (cydney_morgan):

I am sorry!

OpenStudy (abhisar):

It's ok \(\Huge{\overset{\frown}{\normalsize \left( \begin{matrix} \Large\cdot \quad \cdot\\ \cdot\\ \huge \smile \end{matrix} \right)}}\normalsize \\ \;/\quad \;\;\quad \backslash\)

OpenStudy (abhisar):

kk\(\color{blue}{\text{Originally Posted by}}\) @Cydney_Morgan Elements in the same group on the periodic table have the same number of electrons. number of protons. <--- My choice \(\huge \cancel{\bcancel{}}\) number of neutrons. number of valence electrons. \(\color{blue}{\text{End of Quote}}\)

OpenStudy (cydney_morgan):

Oh. :( I'm not really sure what the correct answer is now, But my second choice would be the number of electrons. Is that correct? Sorry :(

OpenStudy (abhisar):

\(\color{blue}{\text{Originally Posted by}}\) @Cydney_Morgan Oh. :( I'm not really sure what the correct answer is now, But my second choice would be the \(\cancel{\bcancel{number of electrons.}}\) Is that correct? Sorry :( \(\color{blue}{\text{End of Quote}}\)

OpenStudy (cydney_morgan):

Darn it! Let me go look up a periodic table.

OpenStudy (cydney_morgan):

OH it's the number of Valence electrons! Is that correct? Source: http://scienceline.ucsb.edu/getkey.php?key=990

OpenStudy (abhisar):

\(\color{blue}{\text{Originally Posted by}}\) @Cydney_Morgan OH it's the number of Valence electrons! Is that correct? \(\huge\checkmark\) Source: http://scienceline.ucsb.edu/getkey.php?key=990 \(\color{blue}{\text{End of Quote}}\)

OpenStudy (abhisar):

For any neutral element No of protons=No of electrons. So if no. of proton is not equal then no of electron can also not be equal.

OpenStudy (cydney_morgan):

All the atomic numbers, protons, are different right? So my first choice was rather dumb then.

OpenStudy (abhisar):

Elements in same group have same number of valence shell electron and thus have similar valency. This is the reason for their very similar chemical properties.

OpenStudy (abhisar):

\(\color{blue}{\text{Originally Posted by}}\) @Cydney_Morgan All the atomic numbers, protons, are different right? So my first choice was rather dumb then. \(\color{blue}{\text{End of Quote}}\) I was about to say that......! (not the dumb part) If protons will be same then atomic number will also be same and then all the elements in the group will be same !

OpenStudy (cydney_morgan):

Oops! I'm extremely glad I checked with you then! If I need to double check anything is it ok for me to tag you? I do not want you to think I am cheating (I am not)

OpenStudy (abhisar):

yeah sure, go ahead ! and i know u r not cheating...cheaters don't interact !

OpenStudy (cydney_morgan):

True, Thank you so much! I'll keep studying!

OpenStudy (abhisar):

\(\huge\sf\color{green}{\text{✌゚\(\ddot\smile\) ✌゚}}\)